OpenStudy (anonymous):

Seriously guys come on help me with this please, I'd do the same for you. Find the equation of the tangent to the graph at the indicated point f(x)= x^2-9; a=2

terenzreignz (terenzreignz):

Aren't you sweet? >:) Tangent to the graph? This screams *derivative*. Can you find the derivative of f(x) ?

OpenStudy (anonymous):

I can solve this problem up to a point then I just get stuck, I am somewhat familiar with derivatives but we are just barely learning them now in class. So I'm doing the best I can with how very little my teacher has gone over.

terenzreignz (terenzreignz):

Okay, so show me what you've got so far, then.

OpenStudy (anonymous):

So far I have gotten to this point 5). 4h+h^2-10/h 1). And I got that far by using this form f(a+h)-f(a)/h 2). (2+h)^2-9-(2^2-9)/h 3). 4+4h+h^2-9-5/h 4). 4+4h+h^2-14/h After step 5 I just get completely lost

terenzreignz (terenzreignz):

I suggest using this form instead: \[\Large f'(a) = \lim_{x\rightarrow a}\frac{f(x)-f(a)}{x-a}\] It might actually be a little easier

OpenStudy (anonymous):

Unfortunately our teacher never taught us that formula only f(a+h)-f(a)/h and then even still I'm generally getting it I just get stuck at the second to last step.

terenzreignz (terenzreignz):

Okay, no problem. \[\Large \lim_{h\rightarrow0}\frac{f(a+h)-f(a)}{h}\]

terenzreignz (terenzreignz):

You get \[\Large \lim_{h\rightarrow0}\frac{\color{blue}{(a+h)^2-9}-\color{red}{(a^2-9)}}{h }\] Please simplify :)

OpenStudy (anonymous):

1). (2+h)^2-9-(2^2-9)/h 2). 4+4h+h^2-9-5/h 3). 4+4h+h^2-14/h 4). 4h+h^2-10/h <------this is where I get stuck not sure what to do after this point

OpenStudy (anonymous):

algebra mistake

OpenStudy (anonymous):

how do you mean?

OpenStudy (anonymous):

if you do it right, using your method, everything without an \(h\) in it should cancel (add up to zero) in the numerator

OpenStudy (anonymous):

Oh I see I have reworked this problem so many times not sure where I went wrong in my math.

OpenStudy (anonymous):

\[f(2)=2^2-9=-5\] \[f(2+h)=(2+h)^2-9=4+4h+t^2-9=-5+4h+4h^2\] \[f(2+h)-f(2)=-5+4h+4h^2-(-5)=4h+h^2\]

OpenStudy (anonymous):

Great thanks for that, I am kind of confused as to how you got what you got. When I do the problem I get something totally different.